# Test-Time Alignment of LLMs via Sampling-Based Optimal Control in pre-logit space

### v1

@@ C.2 Hyper-parameter tuning
We tune the hyperparameter to optimize reward values for randomly selected 10 training data prompts for BoN, ARGS, and AISP. The hyperparameter of RE-Control is tuned on states and reward pairs collected on test dataset following (Kong et al., 2024). When there are multiple hyperparameters, we performed grid search. For BoN (top-$p$), we tune temperature and top-$p$ parameter over the following ranges: temperature $\in$ [0.4, 0.6, 0.8, 1.0] and $p\in[0.7,0.8,0.9,0.95]$. For ARGS, we tune the weight for the reward value $w$ over the range: [1e-05, 1e-04, …, 100, 1000]. Top-k is set to 32, which corresponds to $n$ of AISP. For RE-Control, we tune learning rate of value function over the range: [1e-05, 1e-04, …, 1.0, 10]. The other hyperparameters follow the settings in the code of (Kong et al., 2024) and we use three layer MLP. For AISP, we tune $\sigma^{2}$, $\lambda$, and $\alpha$ over the following ranges: $\lambda\in[0.1,0.3,0.5,0.7]$ for UltraRM and $\lambda\in[60,120,240,480]$ for Eurus, $\sigma^{2}\in[0.1,0.3,0.5,0.7]$, $\alpha\in[0.99,0.999,0.9999,0.99999]$. Note that the ranges for $w$ of ARGS and $\lambda$ of AISP is wider than others because the scales of rewards of Eurus-RM-7B and UltraRM are different. Selected hyper-paramters of AISP is listed in Tab. 4
@@
$\sigma^{2}$ 0.5 0.7 0.5 0.7 0.7 0.7
$\lambda$ 0.3 240 0.1 480 0.5 60
$\alpha$ 0.999 0.999 0.9999 0.99 0.999 0.999
@@ C.3 Hyper-parameters for generations
As reference, we used the code of BoN in trl [^4] https://github.com/huggingface/trl/blob/main/trl/extras/best_of_n_sampler.py for the implementation of baseline method. We referred to the public codes for inference-time alignment to help determine the experimental setup [^5] https://github.com/Lingkai-Kong/RE-Control; https://github.com/deeplearning-wisc/args especially for evaluations of diversity, coherence, and evaluation by GPT-4.
Following (Kong et al., 2024), a begin word ”User:” is added to prompts for Llama3_8B and ”Human” is added for Vicuna_7B, and finish words of ”Assistant:” are added as preprocessing. We also add ”User:” in the prompt for Gemma3_4B.
@@ C.6 Computation of KL divergence
$\mathbb{D}_{\mathrm{KL}}(P_{\mathrm{AISP}}(\mathbf{y}|\mathbf{x})|P_{\mathrm{LLM}}(\mathbf{y}|\mathbf{x}))$ $=P_{\mathrm{AISP}}(\mathbf{y}|\mathbf{x})\log\frac{P_{\mathrm{AISP}}(\mathbf{y}|\mathbf{x})}{P_{\mathrm{LLM}}(\mathbf{y}|\mathbf{x})}$ (59)
@@
$\log\frac{\prod_{t}P_{AISP}(y_{t}|\bm{y}_{<t})}{\prod_{t}P_{LLM}(y_{t}|\bm{y}_{<t})}=\log(\prod_{t}P_{AISP}(y_{t}|\bm{y}_{<t}))-\log\prod_{t}P_{LLM}(y_{t}|\bm{y}_{<t})$ (64)
$=\sum_{t}\left[\bm{w}_{i}^{\top}\bm{u}_{t}^{*}-\log(\sum_{j=1}^{|\mathcal{V}|}\exp(\bm{w}_{j}^{\top}\bm{z}_{t}+\bm{b}_{j}))+\log(\sum_{j=1}^{|\mathcal{V}|}\exp(\bm{w}_{j}^{\top}(\bm{z}_{t}+\bm{u}_{t})+\bm{b}_{j}))\right].$ (65)
@@ D.3 Average rewards for different settings of $\kappa$, $n$ and NN
Appendix E LLM Usage
In addition to using LLMs for experiments, we utilized LLMs to correct grammatical errors and to rephrase some sentences to improve the naturalness of English expressions of some parts of this paper. This procedure was performed by giving the LLMs just partial sentences consisting of few words. We did not perform text generation for larger texts, such as paragraphs exceeding several lines.

### v2

@@ C.2 Hyper-parameter tuning
We tune the hyperparameter to optimize reward values for randomly selected 10 training data prompts for BoN, ARGS, and AISP. The hyperparameter of RE-Control is tuned on states and reward pairs collected on test dataset following (Kong et al., 2024). When there are multiple hyperparameters, we performed grid search. For BoN (top-$p$), we tune temperature and top-$p$ parameter over the following For BoN ($\mathcal{N}$), we tune $\sigma^{2}$ over the range: [0.1, 0.3, 0.5, 0.7, 1.0]. ranges: temperature $\in$ [0.4, 0.6, 0.8, 1.0] and $p\in[0.7,0.8,0.9,0.95]$. For ARGS, we tune the weight for the reward value $w$ over the range: [1e-05, 1e-04, …, 100, 1000]. Top-k is set to 32, which corresponds to $n$ of AISP. For RE-Control, we tune learning rate of value function over the range: [1e-05, 1e-04, …, 1.0, 10]. The other hyperparameters follow the settings in the code of (Kong et al., 2024) and we use three layer MLP. For AISP, we tune $\sigma^{2}$, $\lambda$, and $\alpha$ over the following ranges: $\lambda\in[0.1,0.3,0.5,0.7]$ for UltraRM and $\lambda\in[60,120,240,480]$ for Eurus, $\sigma^{2}\in[0.1,0.3,0.5,0.7]$, $\alpha\in[0.99,0.999,0.9999,0.99999]$. Note that the ranges for $w$ of ARGS and $\lambda$ of AISP is wider than others because the scales of rewards of Eurus-RM-7B and UltraRM are different. Selected hyper-paramters of AISP is listed in Tab. 4
@@
$\sigma^{2}$ 0.7 0.5 0.5 0.5 0.5 0.5
$\lambda$ 0.5 60 0.3 60 0.7 60
$\alpha$ 0.999 0.99999 0.99999 0.9999 0.999 0.999
@@ C.3 Hyper-parameters for generations
As reference, we used the code of BoN in trl [^4] https://github.com/huggingface/trl/blob/main/trl/extras/best_of_n_sampler.py for the implementation of baseline method. We referred to the public codes for inference-time alignment to help determine the experimental setup [^5] https://github.com/Lingkai-Kong/RE-Control; https://github.com/deeplearning-wisc/args especially for evaluations of diversity, coherence, and evaluation by GPT-4. We found that the preprocessing of HHRLHF is not applied in the public code of (Kong et al., 2024) and modified it. This preprocessing is removal of the response of datasets and the following begin word procedure: Following (Kong et al., 2024), a begin word ”User:” is added to prompts for Llama3_8B and ”Human” is added for Vicuna_7B, and finish words of ”Assistant:” are added as preprocessing. We also add ”User:” in the prompt for Gemma3_4B.
@@ C.6 Computation of KL divergence
$\mathbb{D}_{\mathrm{KL}}(P_{\mathrm{AISP}}(\mathbf{y}|\mathbf{x})|P_{\mathrm{LLM}}(\mathbf{y}|\mathbf{x}))$
$=$ $P_{\mathrm{AISP}}(\mathbf{y}|\mathbf{x})\log\frac{P_{\mathrm{AISP}}(\mathbf{y}|\mathbf{x})}{P_{\mathrm{LLM}}(\mathbf{y}|\mathbf{x})}$
@@
$\log$ $\frac{\prod_{t}P_{AISP}(y_{t}|\bm{y}_{<t})}{\prod_{t}P_{LLM}(y_{t}|\bm{y}_{<t})}=\log(\prod_{t}P_{AISP}(y_{t}|\bm{y}_{<t}))$
$~~~~~~~~~~~~~~~~~-\log\prod_{t}P_{LLM}(y_{t}|\bm{y}_{<t})$ (62)
$=\sum_{t}\left[\bm{w}_{i}^{\top}\bm{u}_{t}^{*}-\log(\sum_{j=1}^{|\mathcal{V}|}\exp(\bm{w}_{j}^{\top}\bm{z}_{t}+\bm{b}_{j}))\right$.
$~~~~~~~~~~~\left.+\log(\sum_{j=1}^{|\mathcal{V}|}\exp(\bm{w}_{j}^{\top}(\bm{z}_{t}+\bm{u}_{t})+\bm{b}_{j}))\right].$ (63)
@@ D.3 Average rewards for different settings of $\kappa$, $n$ and NN
D.4 Additional tasks
To investigate the effectiveness of AISP in various tasks, we compare AISP with BoN on Alpaca-Eval (Li et al., 2023), GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and TruthfulQA (Lin et al., 2022). For the last two tasks, we use the lm-evaluation-harness codes (Gao et al., 2024). We limit token length of responses to 128. The hyper-parameters are the same with the evaluation on SHP with Llama3_8B and UltraRM/Eurus: we set $n=32$, $\kappa=32$, and $N=1024$, and the LLM is llama3_8B. We used UltraRM except for GSM8K and used Eurus for GSM8K because the original paper of Eurus (Yuan et al., 2024) have shown effectiveness on GSM8K. Tables 6-9 show that AISP is effective on these tasks.
Table 6: Alpaca-Eval 2.0
Length controlled winrate Win rate Standard error
AISP 5.64 2.86 0.59
BoN 3.95 2.24 0.52
Table 7: GSM8K (8 shot)
Acc
AISP 67.5
BoN 66.0
Table 8: HumanEval
Pass@1
AISP 41.4
BoN 34.1
Table 9: TruthfulQA
BLEU acc ROUGE1 acc
AISP 0.426 0.479
BoN 0.424 0.458